Please state True false or uncertain and explain in 3 sentences why?

1)If an economy produces 2 goods (X and Y) subject to a PPB similar to that found in Fig 1-2 of your text, then the opportunity cost of Xrises as more of it is produced.

2)Suppose that, with an equivalent unit of resources, Canada can produce 6 tons of iron ore or 2 tons of bauxite and Australia can produce 2 tons of iron ore or 1 ton of bauxite. Canada has an absolute advantage in the production of iron while Australia has a comparative advantage in the production of bauxite.

3)The data in 112A1-3 tells us that trading bauxite from Australia to Canada at a price of 2.5 units of iron ore per ton of bauxite will make both countries better off.
Answer #1

1.

False

The opportunity cost will not change when PPF curve will be straight line. Here, marginal rate of substitution, will be constant. As a result, opportunity cost will be fixed.

2.

True

For Canada, opportunity cost of producing bauxite = 6/2 = 3 tons of iron ore

For Australia, opportunity cost of producing bauxite = 2/1 = 2 ton of iron ore

Since Australia has relatively lower opportunity cost in production of bauxite, then Australia has comparative advantage in production of Bauxite. Further, Canada is producing more efficiently of bauxite than Australia. So, Canada has absolute advantage in production of bauxite.

3.

True

Australia can only produce 2 ton of iron ore by foregoing 1 ton of bauxite. So, getting of 2.5 ton of iron ore is in their benefit. Canada has to the opportunity cost of producing iron ore = 2/6 = .33 ton of bauxite

But, trade makes them to get it as 1 ton of iron ore = 1/2.5 = .4 ton of bauxite that is greater than the opportunity cost. So, it is also beneficial for Canada.